Using the information from the table below, calculate and interpret the poverty gap for a country in which 80% of the population is below the poverty line of $1 a day.

Distribution of daily income

Between 0 and 20 cents

Between 20 and 40 cents

Between 40 and 60 cents

Between 60 and 80 cents

Between 80 and 100 cents

Percentage of population

20

15

15

10

20

(b) [ 4 points] What is the advantage of this measure over the headcount ratio?
Answer #1

P = (1/n) * \tiny \sum (L - Ci)/L { i -> 1 to q}

P = (1/100)*(20*(1-0.1)/1 + 15*(1-0.3)/1 + 15*(1-0.5)/1 + 10*(1-0.7)/1 + 20*(1-0.9)/1)

P = (1/100)*(41)

P = 0.41 = 41%

The poverty gap is defined as the extent to which individuals on average fall below the poverty line, and expresses it as a percentage of the poverty line, or, in other words, the average difference between poor households' income and the poverty line. Therefore, the average difference between poor households' income and the poverty line is 41% for this country.

The poverty gap index is an improvement over the poverty measure headcount ratio which simply counts all the people below a poverty line, in a given population, and considers them equally poor. Poverty gap index estimates the depth of poverty by considering how far, on the average, the poor are from that poverty line.